# Conversational Query Reformulation with the Guidance of Retrieved Documents

### v3

@@ 5 Conclusion
We present the GuideCQR, a novel query reformulation framework for conversational search. GuideCQR effectively reformulates conversational queries to be more retriever-friendly by leveraging guided documents. GuideCQR not only better aligns queries with user intent but also significantly enhances both the precision and relevance of retrieval results, offering a meaningful advancement in conversational search systems. Experimental results across various datasets and metrics confirm the capability of GuideCQR over human rewritten query and previous CQR methods.
@@ Appendix A Implementation Details
To re-rank the guided documents as in section 3.1.2, we select the ember-v1 †††https://huggingface.co/llmrails/ember-v1 and mxbai-embed-large-v1 ‡‡‡https://huggingface.co/mixedbread-ai/mxbai-embed-large-v1 models as the first and second re-ranking Sentence-Transformer models, respectively. And we use OpenAI text-embedding-3-small for generating embeddings for keywords and answers in the filtering stage in GuideCQR. We use different embedding models at each stage to mitigate potential biases caused by a single model. We use KeyBERT (Grootendorst, 2020) for augmenting keywords. And we use RoBERTa-Base-Squad2 §§§https://huggingface.co/deepset/roberta-base-squad2 for generating the expected answer. We use the most common query among the five candidate queries provided by the baseline Mao et al. (2023) during the re-ranking, answer generation and filtering process in CAsT datasets. This is because Mao et al. (2023) provides a set of five candidate queries when generating their queries, so we make representative among them. And in augmenting keyword process, the number of keywords extracted can be fewer than the span length because the number of keywords reasonable in each document can be less than the length of the keyword span. This means that the keyword span length is maximum number of keywords for each document.
We conduct our experiments using an AMD EPYC 7313 CPU (3.0 GHz) paired with four NVIDIA RTX 4090 GPUs. We use Python 3.11.5 and PyTorch 2.3.1 for the software environment .
Appendix B Query Analysis
Appendix C Selecting FilterScore
@@
We experiment with various settings to find the optimal hyperparameter with the validation set for the proposed method. Before filtering, we use top-4 span 15 keywords and top-10 answers for CAsT-19, and top-5 span 5 keywords and top-10 answers for CAsT-20 and top-1 span 10 keywords and top-10 answer for QReCC. And GuideCQR uses these keywords and answers as our final query set. We select the for each keyword and answer by experimenting with validation sets for the datasets to find the optimal scores. We achieve our final results using a of (1, 1.9) for each keyword and answer in CAsT-19, (0.1, 1.95) in CAsT-20, and (0.5, 9) in QReCC.
Appendix D Process of generating of CAsT
Appendix E Dialog history usage in GuideCQR
Appendix F Dataset Statistics
@@
Table 8 presents statistics for the CAsT datasets, including the number of conversations, queries, and documents. Both CAsT datasets share the same document collection and provide passage-level relevance judgments, as well as human rewrites for each turn.

### v4

@@ 5 Conclusion
We present GuideCQR, a novel query reformulation framework utilizing initially retrieved documents from query set for conversational search. GuideCQR effectively reformulates conversational queries to be more retriever-friendly by extracting meaningful information from these guided documents. Experimental results across various datasets and metrics confirm the capability of GuideCQR over human rewritten query and previous CQR methods. We also show that GuideCQR can be effectively adapted to various types of re-written queries.
@@ Appendix A Implementation Details
To re-rank the guided documents as in section 3.1.2, we select the ember-v1 222https://huggingface.co/llmrails/ember-v1 and mxbai-embed-large-v1 333https://huggingface.co/mixedbread-ai/mxbai-embed-large-v1 models as the first and second re-ranking Sentence-Transformer models, respectively. And we use OpenAI text-embedding-3-small for generating embeddings for keywords and answers in the filtering stage in GuideCQR. We use different embedding models at each stage to mitigate potential biases caused by a single model. We use KeyBERT (Grootendorst, 2020) for augmenting keywords. And we use RoBERTa-Base-Squad2 444https://huggingface.co/deepset/roberta-base-squad2 for generating the expected answer. We use the most common query among the five candidate queries provided by the baseline Mao et al. (2023) during the re-ranking, answer generation and filtering process in CAsT datasets. This is because Mao et al. (2023) provides a set of five candidate queries when generating their queries, so we make representative among them. And in augmenting keyword process, the number of keywords extracted can be fewer than the span length because the number of keywords reasonable in each document can be less than the length of the keyword span. This means that the keyword span length is maximum number of keywords for each document. We use roberta-large-mnli 555https://huggingface.co/FacebookAI/roberta-large-mnli for measuring entailment score between augmented keywords and . In the query relevance analysis, we ensure a fair comparison by standardizing the number of documents in each set to 1,700 before conducting the experiment. We leverage Faiss index Johnson et al. (2019) for ANCE retriever. We conduct our experiments using an AMD EPYC 7313 CPU (3.0 GHz) paired with four NVIDIA RTX 4090 GPUs. We use Python 3.11.5 and PyTorch 2.3.1 for the software environment.
Selecting FilterScore
@@
We experiment with various settings to find the optimal hyperparameter with the validation set for the proposed method. The number of documents for keywords and expected answers and optimal keyword span length varies by dataset; these values were chosen to improve retrieval results and enhance query quality for the retriever. Therefore, before filtering, we use top-4 span 15 keywords and top-10 answers for CAsT-19, and top-5 span 5 keywords and top-10 answers for CAsT-20 and top-1 span 10 keywords and top-10 answers for QReCC. And GuideCQR uses these keywords and answers as our final query set. We select the for each keyword and answer by experimenting with validation sets for the datasets to find the optimal scores. We achieve our final results using a of (1, 1.9) for each keyword and answer in CAsT-19, (0.1, 1.95) in CAsT-20, and (0.5, 9) in QReCC.
| CAsT-19 | CAsT-20 | QReCC | Avg | |||||||
| Methods | MRR | NDCG@3 | R@10 | MRR | NDCG@3 | R@10 | MRR | NDCG@3 | R@10 | |
| RawQuery | 31.9 | 13.1 | 3.8 | 9.9 | 6.8 | 2.5 | 6.5 | 5.5 | 11.1 | 10.1 |
| 59.9 | 29.5 | 10.3 | 30.5 | 19.0 | 10.7 | 46.9 | 55.0 | 65.5 | 36.4 | |
| LLM4CS | 67.5 | 41.9 | 11.5 | 53.7 | 38.8 | 20.5 | 47.8 | 45.0 | 69.1 | 44.0 |
| GuideCQR (ours) | 75.4 | 49.7 | 14.0 | 49.1 | 34.4 | 16.4 | 51.2 | 58.6 | 69.0 | 46.4 |
| Human Rewrite | 61.3 | 30.9 | 10.9 | 37.7 | 23.9 | 14.2 | 39.7 | 36.2 | 62.5 | 35.3 |
Appendix B Result using Sparse Retrieval
We provide the performance of GuideCQR using sparse retrieval. This result is not included in the main table because prior studies on the CAsT series have predominantly focused on dense retrieval. Consequently, retrieval results for baseline methods are unavailable and the rewritten queries are not provided either. Thus, we conduct experiments with the available resources. We perform retrieval using BM25 Robertson et al. (2009), and indexing is performed via Pyserini Lin et al. (2021). For sparse retrieval in LLM4CS Mao et al. (2023), we concatenate the rewritten query and response of generated LLM4CS final query.
As shown in Table 7, GuideCQR significantly enhances performance metrics compared to the across all datasets. GuideCQR achieves state-of-the-art performance in terms of the average score. Similar to earlier observations, the scores for CAsT-20 are lower due to the unique characteristics of the CAsT datasets.
Appendix C Dataset Statistics
@@
Table 8 presents statistics for the CAsT datasets, including the number of conversations, queries, and documents.
@@
Appendix D Performance among Keyword Span Length
| CAsT-19 | CAsT-20 | |||
| MRR | NDCG@3 | MRR | NDCG@3 | |
| 72.2 | 45.2 | 53.4 | 37.2 | |
| 5 span | 76.9 | 49.9 | 57.7 | 41.5 |
| 10 span | 78.8 | 52.0 | 53.7 | 39.8 |
| 15 span | 81.0 | 52.2 | 54.1 | 40.0 |
| 20 span | 80.1 | 52.0 | 54.6 | 39.8 |
Appendix E Performance for removing duplicated keywords
We conduct experiments by removing duplicate keywords, as non-duplicated keywords may lead to improved retrieval performance. Table 11 shows that duplicated keywords performed slightly better than unique ones, so we keep them duplicated.
| MRR | NDCG@3 | |
| Original | 81.0 | 52.2 |
| Duplicate_removed | 78.3 | 49.9 |
Appendix F Process of generating of CAsT
Appendix G Dialog history usage in GuideCQR
Appendix H Query Sample
Appendix I Samples of Positive Queries on Datasets
We provide positive query sample of each dataset. For CAsT, if the top-ranked (first position) retrieved document has a relevance score of 1 or higher, we classify it as a positive case. For the QReCC dataset, we refer to a retrieval result as a positive case if the gold passage is included in the top-1 result. Table 14, Table 15 and Table 16 shows positive query sample of CAsT-19, CAsT-20, and QReCC dataset.
@@ Appendix I Samples of Positive Queries on Datasets
| CAsT-19 Dataset | |
| Previous Turns |
Query: What is throat cancer?
Response: “” Query: Is it treatable? Response: “” Query: Tell me about lung cancer. Response: “” |
| Original Query | What are its symptoms? |
| Human Rewrite | What are lung cancer’s symptoms? |
| LLM | What are the symptoms of lung cancer? |
| GuideCQR Query | What are the symptoms of lung cancer? signs symptoms lung cancer symptoms lung cancer signs symptoms lung cancer cough signs symptoms lung cancer cough (hemoptysis), breath, wheezing, hoarseness coughing; pain chest, coughing |
| Gold Passage (4) | Signs and symptoms of lung cancer typically occur only when the disease is advanced. Signs and symptoms of lung cancer may include: A new cough that doesn’t go away. Changes in a chronic cough or smoker’s cough. Coughing up blood, even a small amount. Shortness of breath. Chest pain. |
| CAsT-20 Dataset | |
| Previous Turns |
Query: How do you know when your garage door opener is going bad?
Response: Light Socket or Logic Board. If the light is still not coming on, it could be the light socket or the logic board. To determine which one it is, follow the steps below. Unplug the garage door opener, then immediately plug the opener back into the electrical outlet. Listen for the light relay on the logic board to click. If you hear a click, replace the light socket. If you do not hear a click, replace the logic board. There are four causes for the lights on a garage door opener not to come on: 1. Light bulb. 2. Contacts in the light socket. 3. Light socket. Query: Now it stopped working. Why? Response: Garage Door Opener Problems. So, when the garage door opener decides to take a day off, it can leave you stuck outside, probably during a rain or snowstorm. Though they may seem complicated, there really are several things most homeowners can do to diagnose and repair opener failures. And, if you are careful not to damage the door or the seal on the bottom of the door, use a flat shovel or similar tool to chip away at the ice. Once you get the door open, clear any water, ice or snow from the spot on the garage floor where the door rests when closed. |
| Original Query | How much does it cost for someone to fix it? |
| Human Rewrite | How much does it cost for someone to repair a garage door opener? |
| LLM | How much does it typically cost to have a professional fix a garage door opener issue? |
| GuideCQR Query | How much does it typically cost to have a professional fix a garage door opener issue? garage door opener repair cost garage door opener repair cost cost repair broken garage door opener garage door opener repair cost garage door opener repair cost fix garage door opener |
| Gold Passage (4) | On average, homeowners report paying $210 to have a garage door opener repaired by a handyman. Most homeowners pay between $170 and $250 for the cost of such a repair project. The minimum reported cost to repair a broken garage door opener is $50, while the most costly repair may amount to $350. If an electrician is needed to repair faulty wiring, the cost of the repair will increase. The minimum reported cost to repair a broken garage door opener is $50, while the most costly repair may amount to $350. If an electrician is needed to repair faulty wiring, the cost of the repair will increase. |
| QReCC Dataset | |
| Previous Turns |
Query: What are the main breeds of goat?,
Answer: Abaza…Zhongwei |
| Original Query | Tell me about boer goats. |
| Human Rewrite | Tell me about boer goats. |
| LLM | What can you tell me about boer goats in relation to the main breeds of goats? |
| GuideCQR Query | What can you tell me about boer goats in relation to the main breeds of goats? goat boer breed south developed |
| Gold Answer | The Boer goat is a breed of goat that was developed in South Africa in the early 1900s for meat production. Their name is derived from the Afrikaans (Dutch) word boer, meaning farmer. |
| Gold Passage | Boer goat - Wikipedia CentralNotice Boer goat From Wikipedia, the free encyclopedia Jump to navigation Jump to search A commercial Boer goat buck Commercial Boer goat buck Boer goat, Doe The Boer goat is a breed of goat that was developed in South Africa in the early 1900s and is a popular breed for meat production. Their name is derived from the Afrikaans ( Dutch ) word boer , meaning farmer. Contents 1 Origins and characteristics 2 Doe 3 Crossbreeding 4 Show goats 5 References 6 External links Origins and characteristics [ edit ] The Boer goat was probably bred from the indigenous South African goats kept by the Namaqua , San , and Fooku tribes, with some crossing of Indian and European bloodlines being possible. They were selected for meat rather than milk production; due to selective breeding and improvement, the Boer goat has a fast growth rate and excellent carcass qualities, making it one of the most popular breeds of meat goat in the world. Boer goats have a high resistance to disease and adapt well to hot, dry semideserts. United States production is centered in west-central Texas , particularly in and around San Angelo and Menard . The original US breeding stock came from herds located in New Zealand. Only later were they imported directly from Africa Boer goats commonly have white bodies and distinctive brown heads. Some Boer goats can be completely brown or white or paint, which means large spots of a different color are on their bodies. Like the Nubian goat , they possess long, pendulous ears. They are noted for being docile, fast-growing, and having high fertility rates. Does are reported to have superior mothering skills as compared to other breeds. |